How high is a tree that casts a 30-ft shadow at the same time a 6-ft post cast a shadow which is 13-ft long?
Answer by jim_thompson5910(35256)

6/13 = x/30

6*30 = 13x

180 = 13x

180/13 = x

x = 180/13

x = 13.846

So the tree is roughly 13.846 ft high.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------